Solving Exponential Equation Practice:

  1. 23x = 49
  2. 32x = 910
  3. 105x+3 = 100011
1
Expert's answer
2021-05-18T07:05:12-0400

"2^{3x}=4^9"

"2^{3x}=2^{18}"

Hence

"3x=18"

"x=6"



"3^{2x}=9^{10}"

"3^{2x}=3^{20}"

Hence. "2x=20"

"x= 10"



"10^{5x+3}=1000^{11}"

"10^{5x+3}=10^{33}"

Hence. "{5x+3}=33"

"5x=30"

"x=6"